# CaveAgent: Transforming LLMs into Stateful Runtime Operators

### v1

@@ Abstract:
LLM-based agents are increasingly capable of complex task execution, yet current agentic systems remain constrained by text-centric paradigms. Traditional approaches rely on procedural JSON-based function calling, which often struggles with long-horizon tasks due to fragile multi-turn dependencies and context drift. In this paper, we present CaveAgent, a framework that transforms the paradigm from "LLM-as-Text-Generator" to "LLM-as-Runtime-Operator." We introduce a Dual-stream Context Architecture that decouples state management into a lightweight semantic stream for reasoning and a persistent, deterministic Python Runtime stream for execution. In addition to leveraging code generation to efficiently resolve interdependent sub-tasks (e.g., loops, conditionals) in a single step, we introduce Stateful Runtime Management in CaveAgent. Distinct from existing code-based approaches that remain text-bound and lack the support for external object injection and retrieval, CaveAgent injects, manipulates, and retrieves complex Python objects (e.g., DataFrames, database connections) that persist across turns. This persistence mechanism acts as a high-fidelity external memory to eliminate context drift, avoid catastrophic forgetting, while ensuring that processed data flows losslessly to downstream applications. Comprehensive evaluations on Tau2-bench, BFCL and various case studies across representative SOTA LLMs demonstrate CaveAgent’s superiority. Specifically, our framework achieves a 10.5% success rate improvement on retail tasks and reduces total token consumption by 28.4% in multi-turn scenarios. On data-intensive tasks, direct variable storage and retrieval reduces token consumption by 59%, allowing CaveAgent to handle large-scale data that causes context overflow failures in both JSON-based and CodeAct-style agents. Furthermore, the accessible runtime state provides programmatically verifiable feedback, establishing a rigorous foundation for future research in Reinforcement Learning (RL).
= Date: Jan 5, 2026
= Main Contact: Zhenglin Wan (vanzl@u.nus.edu), Jun Song (junsong@hkbu.edu)
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable general knowledge acquisition and human-like reasoning capabilities, exhibiting exceptional performance across diverse natural language processing tasks. Building upon these foundational capabilities, tool-integrated reasoning (TIR) has enhanced LLM agent through reasoning to interact with external tools or application programming interfaces (APIs) in a multi-turn manner [^1] In this paper, we refer to tool-use and function-calling interchangeably. (lu2023chameleon, shen2023hugginggpt, patil2024gorilla, qu2025tool), thereby substantially expanding their information access and solution space. This largely amplifies the landscape of LLM agents to variety of domains, such as scientific discovery (boiko2023emergent, bran2024chemcrow), mathematical problem-solving (gao2023pal, chen2022program), Web GUI navigation (zhou2023webarena, yao2022webshop) and Robotics (driess2023palme, brohan2023rt2).
Despite the promising landscape, the conventional protocol for tool use requires LLMs to conform to predefined JSON schemas and generate structured JSON objects containing precise tool names and arguments (qin2023toolllm, openai2023gpt4). For example, to retrieve stock data, the model must strictly synthesize a JSON string like {"tool": "get_stock", "params": {"ticker": "AAPL", "date": "today"}}, requiring exact adherence to syntax and field constraints. However, this approach exhibits significant limitations: 1) Flexibility: Agents are typically constrained by a rigid, iterative loop: executing a single tool call (or a parallel batch), serializing the output, and feeding the result back into the context for the subsequent generation. This introduces significant latency and context redundancy, resulting in suboptimal performance when addressing complex tasks that demand the sophisticated orchestration of sequential tool interactions (wu2023autogen, shen2023hugginggpt). 2) Hallucination: Achieving reliable tool-calling capabilities necessitates that the LLM outputs tool-related tokens with zero-shot precision (patil2024gorilla). However, in practice, relying on in-context learning to guide tool generation often suffers from severe hallucinations, such as inventing non-existent parameters or violating type constraints (patil2024gorilla, qin2023toolllm, li2023apibank). Crucially, errors in early turns propagate through the conversation, leading to cascading failures in multi-turn tasks (kim2023language). Moreover, adhering to JSON-Schema needs post-training LLMs which requires significant time and computational resources, and may even result in an LLM with lower intelligence levels than before post-training (JSON-post).
@@
While recent works attempt to address this issue by empowering LLMs with code-based tool-use (wang2024executable, yang2024if), they predominantly adopt a process-oriented paradigm where the runtime state remains internalized and text-bound. The interaction forces a "textualization bottleneck”: variables are accessible to external systems only through text output, requiring serializing into text strings (e.g., printing a DataFrame) to communicate with the user (wang2024executable, yao2022react). This limitation fundamentally prohibits the direct input and output of structured, manipulatable objects, making it inefficient or impossible to handle complex non-textual data (e.g., large datasets, videos) (qiao2023taskweaver) and interact with down-stream tasks. To address these limitations:
@@
We present CaveAgent, an open-source platform that pioneers the concept of Stateful Runtime Management in LLM agents. This marks a shift of code-based function calling paradigm from "process-oriented function-calling” to persistent "object-oriented state manipulation”. CaveAgent operates on a dual-stream architecture that enhances interaction between LLM agents and environments through two distinct streams: a semantic stream for reasoning and a runtime stream for state management and code execution. In this framework, the semantic stream remains lightweight, receiving only abstract descriptions of functions’ API and variables. It leverages the LLM’s inherent coding capabilities to generate code that manipulates the runtime stream—the primary locus of our stateful management. By injecting complex data structures (such as graph, dataframe, etc.) directly into the runtime as persistent objects, we achieve another form of context engineering: the agent manipulates high-fidelity data via concise variable references, decoupling storage from the limited context window to the persistent runtime. Specifically, any intermediate result (e.g., DataFrames, planning trees, or key metadata) can be stored in newly injected stateful variables and the agent will actively retrieve relevant variables for later use or downstream applications. This avoids catastrophic forgetting (catastrophic1, catastrophic2), enables efficient context compression and error-free recall for long-term memory where the runtime serves as an "external memory dictionary". Besides, this persistent environment further enables few-steps solution of complex logical dependencies by directly using code to interact with multiple logically inter-dependent tools, allowing the agent to compose intricate workflows (e.g., data filtering followed by analysis) in a few turns, thus avoiding the potential error and instability caused by multi-round function calling (multi-turn1, qin2023toolllm). Furthermore, the runtime’s transparency makes agent behavior fully verifiable, supporting checks on both intermediate programmatic states and final output objects of any data type. This capability creates a rigorous framework for Reinforcement Learning by enabling the generation of verifiable, fine-grained reward signals. Finally, CaveAgent supports lossless artifact handoff by returning native Python objects rather than text representations, and the extraction of manipulated Python objects for direct use in down-stream tasks such as UI rendering, visualization and structured validation. The runtime can be easily serialized and reloaded, providing a simple yet powerful mechanism for preserving the agent’s complete state across sessions and enabling true long-term memory and task continuity. This transforms the LLM from isolated text generator into an interoperable computational entity, seamlessly embedding within complex software ecosystems and automated decision-making frameworks.
@@
In addition to these insights, we found that the function calling paradigm in CaveAgent could potentially extends beyond single-agent capabilities to pioneer Runtime-Mediated Multi-Agent Coordination, as shown in Figure 1 and the right-bottom sub-figure of Figure 2. Unlike conventional frameworks where agents coordinate via lossy, high-latency text message passing (li2023camel, generative_agents), CaveAgent enables agents to interact through direct state manipulation. In this paradigm, a supervisor agent can programmatically inject variables into a sub-agent’s runtime to dynamically alter its environment or task context, effectively controlling behavior without ambiguous natural language instructions. Furthermore, multiple agents can operate on a unified shared runtime, achieving implicit synchronization: when one agent modifies a shared object (e.g., updating a global ’weather’ entity in a town simulation), the change is instantly perceivable by all peers through direct reference. This transforms multi-agent collaboration from a complex web of serialized dialogue into a precise, verifiable state flow, ensuring that large-scale coordination remains coherent and grounded (the details of Runtime-Mediated Multi-Agent Coordination can be found in Appendix E). We summarize our contribution as follows:
@@
  • We introduce CaveAgent, a new function-calling paradigm that pioneers the concept of Stateful Runtime Management. This architecture marks a paradigm shift from process-oriented function calling to persistent, object-oriented state management. CaveAgent achieves a form of context compression and context-grounded memory recall via delegating context engineering to persistent runtime, eliminating the token overhead and precision loss inherent in textual serialization while enabling the few step solution of complex, logically interdependent tasks.
  • The framework’s programmatic inspectability provides deterministic feedback on intermediate states, establishing a rigorous foundation for future research in Reinforcement Learning with Verifiable Rewards (RLVR) on this paradigm without the need for subjective human annotation.
@@
  • We conduct evaluations demonstrating CaveAgent’s tool use ability on standard benchmarks (e.g., Tau2Bench) and provides comprehensive case study across various domains to showcase the unique advantages of CaveAgent. Additionally, we identified the potential to extend the paradigm to enable Stateful Runtime-Mediated Multi-Agent Coordination and provided qualitative results, opening the opportunity for future research on this direction.
2 Background
Figure 3: Evolution of Agentic Tool Use
In this section, we formally formulate the function-calling (tool-use) in LLMs. We consider an LLM agent $\mathcal{M}_{\theta}$ parameterized by $\theta$, tasked with solving a user query $q$. The agent is equipped with a tool library $\mathcal{T}=\{t_{1},t_{2},\dots,t_{N}\}$. Each tool $t_{i}$ is defined by a tuple $(n_{i},d_{i},\rho_{i})$, representing the tool name, description, and parameter space, respectively. The problem is modeled as a multi-step decision process. At time step $k$, given the context history $H_{k}$, the agent generates a response. The history is defined as a sequence of interactions:
$H_{k}=\{q,(r_{1},a_{1},o_{1}),\dots,(r_{k-1},a_{k-1},o_{k-1})\}$ (1)
where $r_{k}$ denotes the internal reasoning (thought), $a_{k}$ denotes the tool action, and $o_{k}$ denotes the execution observation.
ReAct Paradigm
In the traditional ReAct paradigm (yao2022react), both reasoning $r_{k}$ and action $a_{k}$ are generated as contiguous natural language sequences from the model’s vocabulary $\mathcal{V}$. The generation probability is formalized as: $P(r_{k},a_{k}|H_{k})=\prod_{j=1}^{L}P_{\theta}(w_{j}|H_{k},w_{<j})$, where $w$ represents tokens in the sequence $(r_{k},a_{k})$. Crucially, the action $a_{k}$ in ReAct is a raw text string (e.g., "Action: Search[query]") that requires a heuristic parser $\mathcal{P}$ to extract the executable function and parameters. Let $\hat{a}_{k}$ be the raw text output. The effective action is obtained via: $a_{k}=\mathcal{P}(\hat{a}_{k}),\quad\text{s.t.}\quad\hat{a}_{k}\in\mathcal{V}^{*}$. This formulation suffers from hallucination and format errors, as the support of $P_{\theta}$ covers the entire vocabulary $\mathcal{V}$, meaning $\sum_{w\notin\text{valid\_syntax}}P_{\theta}(w|\dots)>0$ (patil2024gorilla, liu2023agentbench).
JSON-Schema Function Calling
To address the ambiguity of unstructured text generation, modern agents adopt JSON-Schema Function Calling (patil2024gorilla, qin2023toolllm). Here, the toolset $\mathcal{T}$ is augmented with a set of structured schemas $\mathcal{S}=\{S_{1},\dots,S_{N}\}$, typically defined in JSON Schema format. The reasoning process remains in natural language, but the action generation is transformed into a constrained decoding process. The model is conditioned explicitly on $\mathcal{S}$, and the action $a_{k}$ is no longer treated as free text, but as a structured object $\mathbf{j}_{k}$ (a JSON object). To help the model output the JSON object correctly, we could utilizes In-Context Learning to internalize schema structures (toolformer). Crucially, special tokens (e.g., <tool>) are introduced to explicitly demarcate the reasoning phase from the action phase. While recent development in Agentic RL adopts Reinforcement Learning using composite reward signal to incentive the model to output correct JSON format and function calling parameter (coderl), this paper mainly focus on inference-time rather than training-time techniques.
Essentially, JSON-Schema function calling operates as a text-based serialization loop (yao2022react, automatedGPT). The process consists of three phases: (1) Context Serialization: The structured schema $\mathcal{S}$ is flattened into a textual description and injected into the system prompt via context engineering; (2) String Generation: The LLM acts as a neural generator, predicting a string JSON payload based on textual instructions; and (3) Execution: An external middleware parses this string, executes the actual code, and serializes the execution result back into text to update the context window. This paradigm does not fundamentally deviate from the traditional context engineering framework of LLMs, suffering from inherent limitations such as context explosion, hallucination, and error propagation (packer2023memgpt, wang2023mint).
Code-based Function Calling
To address these limitations, recent works such as codeact (wang2024executable) utilize executable code as the media of function calling. However, current code agents suffer from architectural limitations. CodeAct essentially does not expose explicit APIs for external object injection and retrieval. Interaction is strictly mediated by the LLM via a "textualization" bottleneck, where intermediate states must be serialized into standard output (e.g., print) to be perceived by the user. For instance, as wang2024executable stated, when the agent requires external data for analysis, Codeact typically downloads the dataset via Python (e.g., pd.read_csv(url)). This approach is inherently inflexible: its interface boundary relies on text serialization for data ingestion, making it difficult to directly inject pre-existing Python objects such as in-memory DataFrames, trained models, or custom class instances without custom workarounds. Moreover, this reliance on text makes it challenging to handle of non-textual or high-dimensionality artifacts (e.g., raw video streams, large-scale databases), and exposes the risk of context explosion, distraction of LLM and hallucination (liu2024lost, packer2023memgpt).
In the subsequent section, we demonstrate how CaveAgent pioneers the Object-Oriented paradigm based on Python’s "everything is an object” philosophy by maintaining two paralleled context stream and delegating context management to a persistent Python runtime stream. Figure 3 shows the evolution path of paradigm shift in Agentic Tool Use.
@@ 3.1 Core Methodologies
In this section, we present the design philosophy of CaveAgent. As illustrated in Figure 4, CaveAgent adopts a dual-stream architecture, maintaining two synchronized streams throughout the interaction lifecycle: a Semantic Stream for light-weight reasoning, and a Runtime Stream for stateful execution, observation and context engineering. This design fundamentally redefines the agentic interaction loop, shifting from stateless text-based serialization to a persistent, state-aware model.
@@
where $h_{t}$ represents the semantic history (we call it "in-prompt context") and $c_{t}$ is the executable code generated by the agent. The critical innovation lies in the decoupling of $h_{t}$ and $\mathcal{S}_{t}$: the semantic stream tracks intent and light-weight reasoning for code generation, while the runtime stream maintains all crucial data and execution state via the code generated by semantic stream.
@@
The core engine of the runtime stream is a persistent Python kernel (specifically, an IPython interactive shell). We conceptualize each interaction turn $t$ not as an isolated API call, but as a cell execution in a virtual Jupyter notebook.
@@
Notably, the runtime stream can assign values to new variables during the interaction process and inject them into the Persistent Namespace (in-runtime context). This enables heavy context in complex tasks, such as large DataFrames, graphs, or other intricate data structures, to be managed entirely by the Python runtime stream as stateful variables. Their values are thus preserved natively in persistent runtime memory without requiring repeated serialization into text, effectively eliminating the risk of hallucination that arises from lossy textual representations. Besides, the agent can inject and store crucial information (such as key reasoning chains and intermediate data analysis results) via new persistent variables into in-runtime context, retaining only a lightweight description and reference in its in-prompt context. Consequently, the runtime functions as an external memory dictionary, allowing the agent to actively retrieve this memory as native, lossless Python objects, thus achieving a form of context compression and avoiding catastrophic forgetting. This property is crucial, as it addresses persistent challenges in agentic tool use—specifically memory, dynamic decision-making, and long-horizon reasoning (bfcl). Meanwhile, this system also makes the manipulation of data objects and propagating them between multiple-turns much easier, no matter how complex the data structures are.
It is also notable that the programmatic state retrieval enables the extraction of manipulated Python objects for direct use in downstream applications. Unlike conventional agents that produce terminal text outputs requiring parsing and reconstruction, CaveAgent exposes native objects (DataFrames, class instances, arrays) with full type fidelity and structural integrity. This enables diverse integration pathways. For example:
  • UI components can bind directly to retrieved objects for exact data visualization, enabling real-time dashboards reflect to exact agent state.
  • RL pipelines can compute precise reward signals through programmatic state inspection rather than noisy text-based heuristics, automate the process of success/failure detection for trajectory labeling, and conduct credit assignment based on state analysis.
  • Validation frameworks can apply unit test assertions and schema verification against returned structures, enabling domain-specific benchmarking with programmatic correctness.
  • Multi-agent systems can pass objects directly between agents without serialization loss, share state synchronization across agent swarm and resolve dependency based on object availability (as discussed in detail at Appendix E).
The agent thus transforms LLMs from an isolated text generator into the operator of a stateful, interoperable computational component whose outputs integrate natively into broader software ecosystems and automated decision-making pipelines.
@@
Parallel to the runtime stream, the semantic stream utilizes LLM as a brain to generate code to manipulate runtime. Besides, it is also responsible for:
@@
  • Observation Shaping: Captures execution outputs and enforces a length constraint $\tau(\cdot)$ to prevent context explosion. This feedback mechanism actively teaches the agent to interact with the persistent state efficiently, prioritizing concise and crucial information over verbose raw dumps in the in-prompt context $h_{t+1}$.
This dual-stream design solves the "Context Explosion" problem: massive data remains in the Runtime Stream ($\mathcal{S}_{t}$), while only the high-level reasoning and necessary summaries flow through the Semantic Stream ($h_{t}$). The LLM effectively operates a remote control (code) to manipulate a complex runtime without needing to hold the runtime’s entire state in its working memory. Compared to traditional JSON-based function calling where larger models tend to parallelize tool calls for efficiency but fall short when there are inter-dependencies between tools (toolsandbox), CaveAgent enables dependency-aware parallelism, allowing agents to dispatch complex, interdependent tool chains in a few turns via executable code to guarantee both efficiency and correctness. Compared to traditional code-based function calling that adopts internalized runtime, CaveAgent opens the runtime as a bidirectional interface, allowing developers to inject arbitrary variables directly and retrieve structured, manipulatable objects of any type at any time and achieves true stateful interoperability. Algorithm 1 in Appendix A showcases the iteration loop of CaveAgent framework. Then, we demonstrate three core designs of CaveAgent beyond traditional JSON-based and code-based tool use.

### v2

@@ Abstract:
LLM-based agents are increasingly capable of complex task execution, yet current agentic systems remain constrained by text-centric paradigms that struggle with long-horizon tasks due to fragile multi-turn dependencies and context drift. We present CaveAgent, a framework that shifts tool use from “LLM-as-Text-Generator” to “LLM-as-Runtime-Operator.” CaveAgent introduces a dual-stream architecture that inverts the conventional paradigm: rather than treating the LLM’s text context as the primary workspace with tools as auxiliary, CaveAgent elevates the persistent Python runtime as the central locus of state, with a lightweight semantic stream serving as its orchestrator. Beyond leveraging code generation to resolve interdependent sub-tasks (e.g., loops, conditionals) in a single step, CaveAgent introduces Stateful Runtime Management: it injects, manipulates, and retrieves complex Python objects (e.g., DataFrames, database connections) that persist across turns, unlike existing code-based approaches that remain text-bound. CaveAgent further provides a runtime-integrated skill management system that extends the Agent Skills open standard, enabling ecosystem interoperability through executable skill injections. This persistence mechanism serves as a high-fidelity external memory that reduces context drift in multi-turn interactions and preserves processed data for downstream applications without information loss. Evaluations on Tau2-bench, BFCL, and case studies across representative SOTA LLMs demonstrate consistent improvements: a 10.5% success rate gain on retail tasks, 28.4% reduction in total token consumption for multi-turn scenarios, and 59% token reduction on data-intensive tasks, enabling CaveAgent to handle data scales that cause context overflow in both JSON-based and code-based agents. The accessible runtime state further provides programmatically verifiable feedback, enabling automated evaluation and reward signal generation without human annotation and establishing a structural foundation for future research in Reinforcement Learning with Verifiable Rewards (RLVR).
[…]
1F4C5 Date: Feb 18, 2026 (v.2)
[…]
1F4BB Code: https://github.com/acodercat/cave-agent
[…]
1F4E7 Main Contact: Zhenglin Wan (vanzl@u.nus.edu), Jun Song (junsong@hkbu.edu.hk)
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated strong knowledge acquisition and reasoning capabilities across diverse natural language processing tasks. Building on these capabilities, tool-integrated reasoning (TIR) enables LLM agents to interact with external tools and APIs in a multi-turn manner [^1] In this paper, we use tool use and function calling interchangeably. (lu2023chameleon, shen2023hugginggpt, patil2024gorilla, qu2025tool), expanding their information access and solution space. This has extended LLM agents to various domains, including scientific discovery (boiko2023emergent, bran2024chemcrow), mathematical problem-solving (gao2023pal, chen2022program), Web GUI navigation (zhou2023webarena, yao2022webshop), and robotics (driess2023palme, brohan2023rt2).
Despite this progress, the conventional protocol for tool use requires LLMs to conform to predefined JSON schemas and generate structured JSON objects containing precise tool names and arguments (qin2023toolllm, openai2023gpt4). For example, to retrieve stock data, the model must synthesize a JSON string like {"tool": "get_stock", "params": {"ticker": "AAPL", "date": "today"}}, requiring exact adherence to syntax and field constraints. This imposes three architectural limitations: 1) Rigid Control Flow: each turn executes a single tool call (or parallel batch) and serializes output back to context, introducing latency for tasks requiring sequential orchestration (wu2023autogen, shen2023hugginggpt); 2) Statelessness: each call is an isolated transaction with no persistent state across turns, so intermediate results must be serialized back into text, causing token overhead for complex data structures and cascading error propagation (qiao2023taskweaver, kim2023language); and 3) Limited Composability: JSON schemas express flat function signatures and cannot natively represent loops, conditionals, or variable dependencies, forcing multi-step logic into error-prone multi-turn dialogue (kim2023language, wang2024executable).
@@
While recent works attempt to address these issues with code-based tool use (wang2024executable, yang2024if), they predominantly adopt a process-oriented paradigm where the runtime state remains internalized and text-bound. This creates a “textualization bottleneck”: variables are accessible to external systems only through text output, requiring serialization into text strings (e.g., printing a DataFrame) to communicate with the user (wang2024executable, yao2022react). This prevents the direct input and output of structured, manipulatable objects, making it inefficient or impossible to handle complex non-textual data (e.g., large datasets, videos) (qiao2023taskweaver) and interact with downstream tasks. To address these limitations:
@@
We present CaveAgent, an open-source framework that introduces the concept of Stateful Runtime Management for LLM agents, shifting code-based tool use from “process-oriented function calling” to persistent “object-oriented state manipulation.” CaveAgent operates on a dual-stream architecture with two distinct streams: a semantic stream for reasoning and a runtime stream for state management and code execution. This represents a fundamental inversion of the conventional paradigm: whereas existing agents treat the LLM’s semantic context as the primary workspace with external tools as auxiliary, CaveAgent elevates the persistent runtime as the primary locus of computation and state, with the semantic stream serving as a lightweight orchestrator that generates code to manipulate it.
By injecting complex data structures (e.g., graphs, DataFrames) directly into the runtime as persistent objects, CaveAgent achieves a form of context engineering: the agent manipulates high-fidelity data via concise variable references, decoupling storage from the limited context window. Any intermediate result (e.g., DataFrames, planning trees, or key metadata) can be stored in persistent variables that the agent actively retrieves for later use or downstream applications (code as action, state as memory). This reduces progressive context degradation in multi-turn interactions (catastrophic1, catastrophic2), enables context compression, and provides error-free recall through the runtime serving as an external memory.
The persistent environment also enables few-step resolution of complex logical dependencies by using code to interact with multiple interdependent tools, allowing the agent to compose workflows (e.g., data filtering followed by analysis) in a few turns rather than through error-prone multi-round function calling (multi-turn1, qin2023toolllm). The runtime’s transparency makes agent behavior fully verifiable, supporting checks on both intermediate programmatic states and final output objects of any data type, thereby enabling fine-grained reward signals for Reinforcement Learning.
Finally, CaveAgent supports artifact handoff without information loss by returning native Python objects rather than text representations, enabling direct use in downstream tasks such as UI rendering, visualization, and structured validation. The runtime can be serialized and reloaded, preserving the agent’s complete state across sessions. This transforms the LLM from an isolated text generator into a stateful, interoperable computational component within broader software ecosystems.
@@
The function-calling paradigm in CaveAgent also extends beyond single-agent capabilities to enable Runtime-Mediated Multi-Agent Coordination, as shown in Figure 1 and Figure 2. Unlike conventional frameworks where agents coordinate via lossy text message passing (li2023camel, generative_agents), CaveAgent enables agents to interact through direct state manipulation. A supervisor agent can programmatically inject variables into a sub-agent’s runtime to alter its environment or task context without ambiguous natural language instructions. Multiple agents can also operate on a unified shared runtime, achieving implicit synchronization: when one agent modifies a shared object (e.g., updating a global “weather” entity in a town simulation), the change is immediately visible to all peers through direct reference. This transforms multi-agent collaboration from serialized dialogue into precise, verifiable state flow (details in Appendix E). We summarize our contributions as follows:
@@
  • We introduce CaveAgent, a tool-use framework built on Stateful Runtime Management. CaveAgent shifts the paradigm from process-oriented function calling to persistent, object-oriented state management. It achieves context compression and context-grounded memory recall by delegating context engineering to a persistent runtime, eliminating the token overhead and precision loss of textual serialization while enabling few-step resolution of logically interdependent tasks. CaveAgent also provides runtime-integrated skill management that extends the Agent Skills open standard with executable injections for ecosystem interoperability.
  • The framework’s programmatic inspectability is a structural property of the architecture: runtime state is deterministically accessible at any point, providing automated evaluation and fine-grained reward signals for Reinforcement Learning with Verifiable Rewards (RLVR) without requiring subjective human annotation.
@@
  • We evaluate CaveAgent on standard benchmarks (e.g., Tau2-Bench) and provide case studies across various domains including geospatial analysis (Appendix G.3) and AutoML multi-agent coordination (Appendix E.4). We also demonstrate the potential to extend the paradigm to Stateful Runtime-Mediated Multi-Agent Coordination with qualitative results, opening directions for future research.
@@ 2.1 Core Methodologies
In this section, we present the design of CaveAgent. As illustrated in Figure 3, CaveAgent adopts a dual-stream architecture, maintaining two synchronized streams throughout the interaction lifecycle: a Semantic Stream for lightweight reasoning, and a Runtime Stream for stateful execution, observation, and context engineering. Unlike conventional agents where the LLM’s text context serves as the primary workspace and external tools are called as auxiliary services, CaveAgent inverts this relationship: the persistent runtime becomes the central locus of data storage, computation, and state management, while the semantic stream is reduced to a lightweight controller that generates code to operate on the runtime.
@@
where $h_{t}$ represents the semantic history (“in-prompt context”) and $c_{t}$ is the executable code generated by the agent. The key design choice is the decoupling of $h_{t}$ and $\mathcal{S}_{t}$: the semantic stream tracks intent and lightweight reasoning for code generation, while the runtime stream maintains all data and execution state via the code generated by the semantic stream.
@@
The execution kernel of the runtime stream is a persistent Python kernel (an IPython interactive shell). We conceptualize each interaction turn $t$ not as an isolated API call, but as a cell execution in a virtual Jupyter notebook.
@@
The runtime stream can also assign values to new variables during interaction and inject them into the persistent namespace (in-runtime context). This enables large context in complex tasks, such as large DataFrames, graphs, or other data structures, to be managed entirely by the Python runtime stream as stateful variables. Their values are preserved natively in persistent runtime memory without requiring repeated serialization into text, eliminating the risk of hallucination from lossy textual representations.
Code as Action, State as Memory  The agent stores key information (such as reasoning chains and intermediate data analysis results) as persistent variables in the runtime context, retaining only a lightweight description and reference in its in-prompt context. The runtime thus functions as an external memory, allowing the agent to retrieve stored data as native Python objects, achieving context compression and reducing progressive context degradation in multi-turn interactions. This property addresses persistent challenges in agentic tool use, specifically memory, dynamic decision-making, and long-horizon reasoning (bfcl).
A natural question arises: what distinguishes storing intermediate results in runtime variables from persisting them to files? We identify three key differences. First, type fidelity: runtime variables preserve native Python object types with full method interfaces (e.g., a DataFrame retains .groupby(), .merge() operations), whereas file-based storage requires serialization that may lose type information or fail entirely for non-serializable objects such as database connections, trained models with custom layers, or open file handles. Second, access latency: runtime variables enable zero-cost retrieval within the same memory space, while file I/O introduces disk latency and requires explicit read/write operations in generated code. Third, lifecycle management: runtime variables are automatically scoped to the agent session and garbage-collected appropriately, whereas file-based approaches require explicit cleanup logic to avoid accumulating temporary artifacts. That said, file-based persistence offers advantages for large-scale data exceeding memory capacity and for checkpointing across agent restarts. We note that CaveAgent also supports storing intermediate results in files, but storing them in runtime variables better respects the persistent runtime paradigm, yielding a more unified system.
Programmatic state retrieval enables the extraction of manipulated Python objects for direct use in downstream applications. Unlike conventional agents that produce text outputs requiring parsing and reconstruction, CaveAgent exposes native objects (DataFrames, class instances, arrays) with full type fidelity. This enables UI rendering via direct object binding, RL reward computation through programmatic state inspection, validation via unit test assertions against returned structures, and lossless object passing in multi-agent systems (Appendix E). The agent thus transforms the LLM from an isolated text generator into the operator of a stateful, interoperable computational component.
@@
Parallel to the runtime stream, the semantic stream uses the LLM to generate code that manipulates the runtime. It is also responsible for:
@@
  • Observation Shaping: Captures execution outputs and enforces a length constraint $\tau(\cdot)$ to prevent context explosion. This feedback mechanism guides the agent to interact with the persistent state efficiently, prioritizing concise and relevant information over verbose raw dumps in the in-prompt context $h_{t+1}$.
This dual-stream design addresses the “Context Explosion” problem: large data remains in the Runtime Stream ($\mathcal{S}_{t}$), while only high-level reasoning flows through the Semantic Stream ($h_{t}$). By making the runtime (not the LLM’s context window) the authoritative store of all data and intermediate state, CaveAgent avoids the token overhead and precision loss inherent in text-centric architectures. Unlike JSON-based function calling, which fails on inter-dependent tool calls (toolsandbox), CaveAgent dispatches interdependent tool chains in a few turns via executable code. Unlike code-based approaches with internalized runtimes, CaveAgent opens the runtime as a bidirectional interface for injecting and retrieving structured objects of any type. Algorithm 1 in Appendix A shows the iteration loop. We next describe the core designs of CaveAgent.